Homework: Homework # 2 Score: 0 of 16 pts Instructor-created question Save 6 of 11 (8 complete) ▼ HW Score: 64%, 64 of 100 p Question Help A friend of yours bought a new sports car with a $5,000 down payment plus a $25,000 car loan that is financed at an interest rate of 0 25% per month for 60 months After 3 years, the Blue Book value of her vehicle in the used-car marketplace is $9,000 a. Calculate the required monthly loan payment on the car b. How much does your friend still owe on the car loan immediately after she makes her 36th payment? c. Compare your answer to Part (b) to $9,000 This situation is called being upside down. What can she do about it? d. If she decides to keep the car but pay $200 more each month, how many months will it take her to payoff the remaining loan she owes? a. The required monthly payment is s (Round to the nearest cent) b. Your friend still owes son the car loan (Round to the nearest dollar,) c. What is the best thing to do in this situation? Choose the correct answer below O A. She should keep the car longer and hope the vehicle remains in good working order O B. In case the mileage is high, the car is worth more than $9,000 and she should consider selling it. ° C. She should sell the car immediately 0 D. She should stop paying out her loan as it is obviously unfair d. It will take her months to payoff the remaining loan balance (Round-up to the nearest month)
Answer #1

1) The loan amount is $25000 and the interest rate is 0.25% for 60 months.

A=P 1+r-1
25000 * ( ( 0.0025*(1.0025)^60) / ( 1.0025)^60 - 1 )
= 449.22

2) Balance after 36th payment

FV = PV(1 + r)n-P FV Future Value(Remaining Balance) PV Present Value(Original Balance) P - Payment r-rate per payment n number of payments

25000 * (1.0025) ^ 36 - 449.22 * ( (1.0025 ^ 36) - 1) / 0.0025
= 10451.38

3) Option A
She should kept the car longer and hope that vehicle remains in good condition.

Obviously, the price of a car is a function of its condition.

4) If she decides to pay $200 more then her EMI would be
449.22 + 200 = 649.22

=NPER(0.25%,649.22,-10451,,1)

= 16.4 months

It will take her 16 months to payoff.